OpenStudy (anonymous):

simplify: (x + 6)(x - 6)

OpenStudy (anonymous):

\[x^2-36\]

OpenStudy (anonymous):

x^2-36

OpenStudy (anonymous):

x^2 -6x + 6x -36 x^2 - 36

OpenStudy (diyadiya):

\[(a+b)(a-b)=a^2-b^2\] \[(x+6)(x-6)=x^2-(6)^2=x^2-36\]